What is the speed of an electron whose Debroglie wavelength is equal to its compton wavelength?


1
Expert's answer
2020-09-06T17:19:34-0400

Solution

Let rest mass of electron"=m_0"

Let Speed of election =v

Then moving mass of electron can be expressed as

"m=\frac{m_0}{\sqrt{1-\frac{v^2}{c^2}}}" . ... eq.1

Where c is speed of light.

De-broglie wavelength is given as

"\lambda=\frac{h}{p}=\frac{h}{mv}" ............. eq. 2

Compton wavelength for electron is


"\lambda_c=\frac{h}{m_0c}" ............ eq. 3

Now according to the question de-broglie wavelength is equal to the compton wavelength for electron

therefore using equations 1,2,3 we got as below

"\frac{h}{mv}=\frac{h}{m_0c}"


"\frac{h\sqrt{1-\frac{v^2}{c^2}}}{m_0v}=\frac{h}{m_0c}"


"\sqrt{1-\frac{v^2}{c^2}}=\frac{v}{c}"

taking square at both sides we get as


"{1-\frac{v^2}{c^2}}=\frac{v^2}{c^2}"

"\frac{2v^2}{c^2}=1"


"\u200b\t\n \\fcolorbox{red}{yellow}{$v=\\frac{c}{\\sqrt{2}}$}"

Therefore speed of electron is "\frac{c}{\sqrt{2}}" .